Find an equation of the parabola whose graph is shown. Square has area 9
Answer #1

Solution

In the given diagram there is a square which has area 9 i.e. if we let a side of square is a then

a^2=9

a=\sqrt{9}

a=3

Now we can see that the vertex of parabola is at origin and directrix passes through the point (-3,0) (dictance of directrix is 3 unit from origin in the left side) and parellel to y-axis then equation of directrix is

x=-3 \text{ or } x+3=0

Now by standard formula as given in following diagram

Focus of given parabola will be at (3,0)

Now the equation of parabola-

y^2=4ax

y^2=4(3)x

y^2=12x